OpenStudy (anonymous):

Eliza started her savings account with $100. Each month she deposits $25 into her account. Write the function that describes the relationship between the amount Eliza deposits and the total amount in her account. Then, determine the average rate of change in the account from the 2nd month to the 10th month.

OpenStudy (anonymous):

@iGreen

OpenStudy (anonymous):

@DEEISME

OpenStudy (igreen):

Started with 100. (0, 100) Add 25: (1, 125) Add 25: (2, 150) Keep adding twenty-five till you get: (10, 350)

OpenStudy (anonymous):

175

OpenStudy (igreen):

Average rate of change formula: \(m = \dfrac{y_2-y_1}{x_2-x_1}\) \(y_2 = 350\) \(y_1 = 150\) \(x_2 = 10\) \(x_1 = 2\) \(m = \dfrac{350-150}{10-2}\) \(m = \dfrac{200}{8}\) \(m = 25\)

OpenStudy (anonymous):

Thanks bro

OpenStudy (anonymous):

The first term of an arithmetic sequence is -5, and the tenth term is 13. Find the common difference.

OpenStudy (anonymous):

8/9 1.8 2 answer choices above

OpenStudy (anonymous):

@CandyCoatedRainDrops

OpenStudy (anonymous):

25x+100

OpenStudy (anonymous):

whoa the answer choices are above